OpenStudy (anonymous):

can someone plz help me?? Evaluate the limit. lim xe^(7/x) − x x →infinity

OpenStudy (inkyvoyd):

Try wolfram alpha

OpenStudy (inkyvoyd):

http://www.wolframalpha.com/input/?i=limit

OpenStudy (dumbcow):

use substitution, or change variable and limit \[\text {Let} \space \space t = e^{7/x} \rightarrow x = \frac{7}{\ln t}\] problem changes to \[\lim_{t \rightarrow 1}\frac{7t}{\ln t}-\frac{7}{\ln t}\] \[=\lim_{t \rightarrow 1}\frac{7t-7}{\ln t} = \frac{0}{0}\] use L'hopitals rule, differentiate top and bottom \[\rightarrow \lim_{t \rightarrow 1}\frac{7}{1/t} = 7\]

OpenStudy (anonymous):

here's another way you might consider: